OpenStudy (anonymous):

y = 3x-7 y = -2x+4 The system above has: a. many solutions b. one solution c. not enough information d. no solutions

OpenStudy (jlg030597):

whats your guess

OpenStudy (jlg030597):

I know the answer but want to see what you think the answer is

OpenStudy (anonymous):

@arcangel72us .. this is what we have in hand y = 3x-7 y = -2x+4 now compare the 2 different values of y i mean 3x - 7 = -2x + 4 try solving it further :D

OpenStudy (anonymous):

@arcangel72us ?? there ??

OpenStudy (anonymous):

well.. i have to leave .... now.....! and am solving the question..! Just ask me out if still any doubt remains... 3x - 7 = -2x + 4 adding +2x to both the sides => 5x -7 = 4 now lets add +7 to both the sides => 5x = 11 => x = 11/5 and since only single value of x is there it means that only 1 value for y exists => only one solution exists :D